OpenStudy (anonymous):

Sketch the graph: y=|x+1|-2 Just need some guidance! Thanks!

OpenStudy (anonymous):

Ok, well lets start with what you know.. How would you graph y=|x| ?

OpenStudy (anonymous):

y=|x+1|-2 To find the x-intercept, substitute in 0 for y and solve for x. (0)=|x+1|-2 Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. |x+1|-2=(0) Remove the parentheses around the expression 0. |x+1|-2=0 Since -2 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 2 to both sides. |x+1|=2 Remove the absolute value term. This creates a \ on the right-hand side of the equation because |x|=\x. x+1=\(2) Set up the + portion of the \ solution. x+1=2 Move all terms not containing x to the right-hand side of the equation. x=1 Set up the - portion of the \ solution. x+1=-(2) Multiply -1 by the 2 inside the parentheses. x+1=-2 Since 1 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 1 from both sides. x=-1-2 Subtract 2 from -1 to get -3. x=-3 The solution to the equation includes both the positive and negative portions of the solution. x=1,-3 To find the y-intercept, substitute in 0 for x and solve for y. y=|(0)+1|-2 Remove the parentheses around the expression 0. y=|0+1|-2 Combine all similar expressions. y=|1|-2 The absolute value of 1 is 1, so the absolute value symbols can be removed. y=1-2 Solve the equation. y=-1 These are the x and y intercepts of the equation y=|x+1|-2. x=1,-3, y=-1